# DiffIR2VR-Zero: Zero-Shot Video Restoration with Diffusion-based Image Restoration Models

### v2

@@ 3.4 Scheduling
SPMCS | PSNR | 20.516 | 21.910 | 20.573 | 20.636 | 21.534 | 21.843 |
SSIM | 0.471 | 0.617 | 0.490 | 0.517 | 0.544 | 0.572 | |
LPIPS | 0.352 | 0.230 | 0.298 | 0.286 | 0.261 | 0.258 | |
0.531 | 0.157 | 1.058 | 8.290 | 0.571 | 0.571 | ||
10.102 | 3.271 | 13.817 | 11.961 | 10.257 | 9.712 | ||
LPIPS | 0.218 | 0.015 | 0.241 | 0.226 | 0.177 | 0.158 | |
DAVIS 4 | PSNR | 23.948 | 25.215 | 23.504 | 23.843 | 23.780 | 24.182 |
SSIM | 0.608 | 0.727 | 0.584 | 0.618 | 0.601 | 0.621 | |
LPIPS | 0.298 | 0.347 | 0.277 | 0.272 | 0.264 | 0.262 | |
0.512 | 0.186 | 0.912 | 0.745 | 0.654 | 0.474 | ||
14.615 | 11.558 | 18.125 | 17.431 | 16.529 | 14.666 | ||
LPIPS | 0.278 | 0.078 | 0.292 | 0.274 | 0.266 | 0.232 | |
DAVIS 8 | PSNR | 22.570 | 22.690 | 20.268 | 20.519 | 21.964 | 22.331 |
SSIM | 0.527 | 0.594 | 0.446 | 0.424 | 0.502 | 0.519 | |
LPIPS | 0.454 | 0.528 | 0.470 | 0.434 | 0.362 | 0.367 | |
0.523 | 0.351 | 2.199 | 1.759 | 0.964 | 0.699 | ||
14.117 | 13.978 | 24.496 | 21.746 | 17.981 | 15.853 | ||
LPIPS | 0.379 | 0.132 | 0.457 | 0.442 | 0.372 | 0.333 |
DiffBIR | ||||||
Metrics | VRT | Shift-Net | VidToMe | Frame | Ours | |
75 | PSNR | 25.050 | 21.033 | 23.791 | 24.585 | 24.520 |
SSIM | 0.787 | 0.381 | 0.618 | 0.649 | 0.649 | |
LPIPS | 0.275 | 0.735 | 0.296 | 0.276 | 0.275 | |
0.314 | 1.757 | 0.765 | 0.751 | 0.706 | ||
17.825 | 27.094 | 21.751 | 21.798 | 21.166 | ||
LPIPS | 0.095 | 0.501 | 0.287 | 0.275 | 0.264 | |
100 | PSNR | 24.582 | 22.573 | 24.606 | 24.524 | 24.534 |
SSIM | 0.744 | 0.484 | 0.676 | 0.648 | 0.652 | |
LPIPS | 0.346 | 0.518 | 0.318 | 0.275 | 0.271 | |
0.294 | 1.126 | 0.781 | 0.763 | 0.696 | ||
17.079 | 23.424 | 21.460 | 21.835 | 20.639 | ||
LPIPS | 0.095 | 0.375 | 0.278 | 0.281 | 0.267 | |
random | PSNR | 24.989 | 21.113 | 23.692 | 24.579 | 24.508 |
SSIM | 0.780 | 0.386 | 0.615 | 0.650 | 0.649 | |
LPIPS | 0.284 | 0.728 | 0.303 | 0.276 | 0.270 | |
0.363 | 1.896 | 0.772 | 0.755 | 0.713 | ||
18.147 | 27.565 | 21.929 | 21.743 | 21.140 | ||
LPIPS | 0.099 | 0.542 | 0.291 | 0.282 | 0.272 |
@@ 4 Experiments
For video super-resolution, we evaluate on SPMCS [95] and DAVIS [59] testing sets, with two downsample scales (4, 8), following the same degradation pipeline of RealBasicVSR [10] to generate LQ-HQ video pairs. For video denoising, we evaluate on REDS30 [54] with 3 different noise levels (std. 75, 100, and std. is uniformly sampled from the range [50, 100]).
@@
We evaluate the restoration performance based on two aspects: (1) image quality, using LPIPS, SSIM, and PSNR; (2) temporal consistency, using warping error , interpolation error, and interpolation LPIPS. Since LPIPS better reflects visual quality, we propose interpolation LPIPS, based on the interpolation error used in a previous study [41], to more accurately measure video continuity from a visual perspective. This involves interpolating a target frame from its previous and next frames and computing the LPIPS between the estimated and target frames.
@@
The experiment is conducted on an NVIDIA RTX 4090 GPU. We apply our method to DiffBIR [85] and SDx4 upscaler [1], both image-based diffusion models, to demonstrate the propsed method’s compatibility with different models. Noted that for models that are restricted to a super-resolution scale of 4, we will apply the process twice and then use bicubic downsampling to achieve 8 results.
@@ 4.1 Comparisons with State-of-the-Art Methods
To verify the effectiveness of our approach, we compare it with several state-of-the-art methods, including FMA-Net [96] for video super-resolution, and VRT [43] and Shift-Net [39] for video denoising. We also compare our method to per-frame restoration and the application of VidToMe [41], a zero-shot video editing method, onto the same model as ours.
@@
As shown in Tab. 1, regression-based methods like FMA-Net [96] perform better on datasets like SPMCS that have minimal motion. However, their generalization ability diminishes significantly with increased motion or severe degradation. VidToMe [41] can generate highly consistent results, but they are often very blurry, leading to poor visual quality. In contrast, our method enhances temporal consistency while maintaining the generation quality of the original diffusion model, making it the most competitive approach. Fig. 7 provides visualizations of two challenging VSR cases. FMA-Net fails to produce sharp results due to domain gaps between training and testing. Diffusion-based image restoration method DiffBIR [45] and SD4 upscaler [1] can generate sharp results with details, while per-frame processing makes the result video temporal inconsistent and jitters across frames. On the contrary, our zero-shot video restoration framework restores a low-quality input video into a temporally consistent high-quality video.
@@
Video denoising, compared to VSR, is a simpler task for regression models, as they can often find the correct pixel value given a sufficiently large batch size. However, our method consistently outperforms others in terms of visual quality (LPIPS) and remains highly robust even as degradation becomes severe. Fig. 8 visualizes the denoising results on the REDS30 dataset. Shift-Net [39] fails to remove all noise, likely due to the out-of-domain noise level; VRT [43] produces smooth results but lacks fine details. Although DiffBIR [45] generates highly detailed images, it suffers from poor temporal consistency, as evident in the changes to the pedestrian’s head and the statue’s face. In contrast, our method preserves both fine details and temporal consistency, effectively balancing these two aspects.
@@ 4.2 Ablation Study
Tab. 2 presents an ablation study comparing different approaches (optical flow and cosine similarity) for finding correspondences and their order in the UNet. As detailed in Sec. 3.3, the hybrid approach of using optical flow at the downsample blocks and cosine similarity at the upsample blocks achieves the best performance. Additionally, our proposed spatial-aware token merging further enhances performance by utilizing spatial information to guide correspondences. The comparisons in Fig. 6 also indicate that our results are smoother, demonstrating better temporal stability.
@@ 5 Conclusion
Our zero-shot video restoration framework has some limitations. Random keyframe sampling may not always select the most representative frames, potentially affecting restoration quality, especially if frames with severe degradation are chosen. Additionally, the sensitivity of the LDM decoder to minor variations in input latents can cause flickering, particularly in dynamic scenes. Future improvements will focus on refining keyframe selection and stabilizing the decoder output to enhance the practical application of diffusion-based video restoration methods.
@@ Appendix A Appendix / supplemental material
A.1 Correspondences identified by cosine similarity without padding removal
Fig. 9 shows that padding value will affect the matching severely.
A.2 Additional Application: Consistent Video Depth
Our zero-shot framework is applicable to any pre-trained image-based diffusion models and could improve the predicted video consistency. Therefore, we integrate our proposed zero-shot framework into a state-of-the-art latent diffusion-based monocular depth estimator: Marigold [35]. Fig. 10 shows that integrating our proposed framework into Marigold helps improve the temporal consistency of video depth estimation.

### v3

@@ 3.4 Scheduling
| DAVIS 4 | PSNR | 23.014 | 25.215 | 23.504 | 23.843 | 23.780 | 24.182 |
| SSIM | 0.566 | 0.727 | 0.584 | 0.618 | 0.601 | 0.621 | |
| LPIPS | 0.405 | 0.347 | 0.277 | 0.272 | 0.264 | 0.262 | |
| 0.520 | 0.186 | 0.912 | 0.745 | 0.654 | 0.474 | ||
| 13.676 | 11.558 | 18.125 | 17.431 | 16.529 | 14.666 | ||
| LPIPS | 0.329 | 0.078 | 0.292 | 0.274 | 0.266 | 0.232 | |
| DAVIS 8 | PSNR | 22.097 | 22.690 | 20.268 | 20.519 | 21.964 | 22.331 |
| SSIM | 0.513 | 0.594 | 0.446 | 0.424 | 0.502 | 0.519 | |
| LPIPS | 0.554 | 0.528 | 0.470 | 0.434 | 0.362 | 0.367 | |
| 0.440 | 0.351 | 2.199 | 1.759 | 0.964 | 0.699 | ||
| 12.624 | 13.978 | 24.496 | 21.746 | 17.981 | 15.853 | ||
| LPIPS | 0.388 | 0.132 | 0.457 | 0.442 | 0.372 | 0.333 | |
| REDS4 4 | PSNR | 23.134 | 25.829 | 24.189 | 24.226 | 24.679 | 25.118 |
| SSIM | 0.589 | 0.761 | 0.638 | 0.641 | 0.657 | 0.683 | |
| LPIPS | 0.357 | 0.327 | 0.247 | 0.242 | 0.211 | 0.222 | |
| 0.579 | 0.392 | 0.817 | 0.811 | 0.704 | 0.499 | ||
| 17.869 | 19.014 | 22.906 | 22.889 | 22.305 | 20.130 | ||
| LPIPS | 0.356 | 0.133 | 0.295 | 0.281 | 0.271 | 0.221 | |
| REDS4 8 | PSNR | 21.894 | 22.842 | - | - | 22.479 | 22.961 |
| SSIM | 0.532 | 0.644 | - | - | 0.559 | 0.59 | |
| LPIPS | 0.538 | 0.423 | - | - | 0.311 | 0.306 | |
| 0.423 | 0.753 | - | - | 0.828 | 0.551 | ||
| 15.502 | 21.519 | - | - | 21.76 | 19.382 | ||
| LPIPS | 0.412 | 0.159 | - | - | 0.351 | 0.287 | |
| REDS4 16 | PSNR | 20.520 | 21.569 | 18.706 | 18.858 | 20.124 | 20.712 |
| SSIM | 0.483 | 0.570 | 0.461 | 0.410 | 0.461 | 0.509 | |
| LPIPS | 0.697 | 0.565 | 0.612 | 0.562 | 0.446 | 0.438 | |
| 0.296 | 0.619 | 2.664 | 2.030 | 1.168 | 0.665 | ||
| 12.945 | 18.758 | 28.478 | 24.000 | 21.33 | 17.731 | ||
| LPIPS | 0.417 | 0.139 | 0.559 | 0.493 | 0.444 | 0.358 | |
| Vid4 4 | PSNR | 19.622 | 23.209 | 20.047 | 20.134 | 20.687 | 21.226 |
| SSIM | 0.425 | 0.679 | 0.478 | 0.473 | 0.497 | 0.525 | |
| LPIPS | 0.491 | 0.375 | 0.343 | 0.331 | 0.329 | 0.326 | |
| 0.687 | 0.203 | 1.502 | 1.397 | 1.156 | 0.677 | ||
| 11.754 | 4.442 | 17.234 | 16.921 | 15.478 | 11.316 | ||
| LPIPS | 0.337 | 0.026 | 0.275 | 0.271 | 0.265 | 0.198 | |
| Vid4 8 | PSNR | 18.811 | 21.033 | 17.813 | 17.992 | 18.636 | 19.304 |
| SSIM | 0.372 | 0.521 | 0.345 | 0.307 | 0.367 | 0.406 | |
| LPIPS | 0.654 | 0.514 | 0.507 | 0.484 | 0.440 | 0.435 | |
| 0.477 | 0.221 | 2.523 | 1.972 | 1.524 | 0.767 | ||
| 9.942 | 5.269 | 22.881 | 19.970 | 18.112 | 12.281 | ||
| LPIPS | 0.393 | 0.032 | 0.423 | 0.419 | 0.395 | 0.294 | |
| DiffBIR | |||||
| Metrics | VidToMe | Shift-Net | Frame | Ours | |
| REDS30 75 | PSNR | 22.671 | 21.033 | 24.585 | 24.520 |
| SSIM | 0.559 | 0.381 | 0.649 | 0.649 | |
| LPIPS | 0.397 | 0.735 | 0.276 | 0.275 | |
| 0.727 | 0.765 | 0.751 | 0.706 | ||
| 18.440 | 21.751 | 21.798 | 21.166 | ||
| LPIPS | 0.375 | 0.501 | 0.275 | 0.264 | |
| REDS30 100 | PSNR | 22.588 | 22.573 | 24.524 | 24.534 |
| SSIM | 0.557 | 0.484 | 0.648 | 0.652 | |
| LPIPS | 0.404 | 0.518 | 0.275 | 0.271 | |
| 0.733 | 1.126 | 0.763 | 0.696 | ||
| 18.370 | 23.424 | 21.835 | 20.639 | ||
| LPIPS | 0.380 | 0.375 | 0.281 | 0.267 | |
| REDS30 random | PSNR | 22.348 | 21.113 | 24.579 | 24.508 |
| SSIM | 0.546 | 0.386 | 0.650 | 0.649 | |
| LPIPS | 0.429 | 0.728 | 0.276 | 0.270 | |
| 0.681 | 1.896 | 0.755 | 0.713 | ||
| 17.608 | 27.565 | 21.743 | 21.140 | ||
| LPIPS | 0.384 | 0.542 | 0.282 | 0.272 | |
| Set8 50 | PSNR | 21.531 | 23.433 | 23.197 | 23.713 |
| SSIM | 0.501 | 0.482 | 0.594 | 0.63 | |
| LPIPS | 0.415 | 0.574 | 0.261 | 0.245 | |
| 0.911 | 1.358 | 1.078 | 0.747 | ||
| 17.217 | 19.845 | 19.732 | 16.814 | ||
| LPIPS | 0.406 | 0.432 | 0.332 | 0.255 | |
| Set8 100 | PSNR | 21.226 | 18.198 | 22.519 | 22.955 |
| SSIM | 0.484 | 0.281 | 0.553 | 0.591 | |
| LPIPS | 0.472 | 0.733 | 0.338 | 0.323 | |
| 0.918 | 2.229 | 1.13 | 0.802 | ||
| 17.367 | 24.661 | 20.18 | 17.444 | ||
| LPIPS | 0.421 | 0.619 | 0.372 | 0.286 | |
| Set8 150 | PSNR | 20.209 | 16.136 | 21.005 | 21.418 |
| SSIM | 0.443 | 0.291 | 0.486 | 0.544 | |
| LPIPS | 0.554 | 0.729 | 0.449 | 0.402 | |
| 0.972 | 4.279 | 1.207 | 0.832 | ||
| 17.872 | 22.343 | 20.729 | 17.616 | ||
| LPIPS | 0.470 | 0.646 | 0.450 | 0.331 | |
@@ 4 Experiments
For video super-resolution, we evaluate on REDS4 (Nah et al., 2019), Vid4 (Liu & Sun, 2013) and DAVIS (Perazzi et al., 2016a) testing sets, with downsample scales 4 and 8, following the degradation pipeline of RealBasicVSR (Chan et al., 2022). For video denoising, we evaluate on REDS30 (Nah et al., 2019) and Set8 (Tassano et al., 2020) with different noise levels (std. 50, 75, 100, 150 and randomly sampled from the range [50, 100]).
@@
We assess (1) image quality via LPIPS, SSIM, and PSNR; (2) temporal consistency, using warping error , interpolation error, and interpolation LPIPS. Since LPIPS better reflects visual quality, we propose interpolation LPIPS, based on the interpolation error used in a previous study (Li et al., 2024), to more accurately measure video continuity from a visual perspective. This involves interpolating a target frame from its previous and next frames and computing the LPIPS between the estimated and target frames.
@@
The experiment is conducted on an NVIDIA RTX 4090 GPU. We apply our method to DiffBIR (Lin et al., 2024) and SDx4 upscaler (sdx, 2023), both image-based diffusion models, to demonstrate the proposed method’s compatibility with different models. Note that for models that are restricted to a super-resolution scale of 4, we will apply the process twice and then use bicubic downsampling to achieve 8 results. However, this will can lead to out-of-memory issues for SDx4 upscaler in REDS.
@@ 4.1 Comparisons with State-of-the-Art Methods
To verify the effectiveness of our approach, we compare it with several state-of-the-art methods, including FMA-Net (Youk et al., 2024) for video super-resolution, and Shift-Net (Li et al., 2023) for video denoising. We also compare our method to per-frame restoration and the application of VidToMe (Li et al., 2024), a zero-shot video editing method, onto the same model as ours. We also try to compare with Upscale-A-Video (Zhou et al., 2023), which is a diffusion-based video restoration model fine-tuned from an image-based diffusion model. However, due to the lack of released implementation, pre-trained weights, or inference results, we cannot include Upscale-A-Video as a competing method in our comparison.
Our zero-shot video restoration framework is designed to be highly adaptable and capable of leveraging a wide range of pre-trained image diffusion models. This flexibility allows easy adaptation from image to video models without extensive retraining, enabling the application of various restoration tasks by simply switching the underlying image diffusion model.
@@
As shown in Tab. 1, regression-based methods like FMA-Net (Youk et al., 2024) struggle with large motion or severe degradation. VidToMe (Li et al., 2024) can generate highly consistent results, but they are often very blurry, leading to poor visual quality. In contrast, our method enhances temporal consistency while maintaining the generation quality of the original diffusion model, making it the most competitive approach. Fig. 6 provides visualizations of two challenging VSR cases. FMA-Net fails to produce sharp results due to domain gaps between training and testing. Diffusion-based image restoration method DiffBIR (Lin et al., 2024) and SD4 upscaler (sdx, 2023) can generate sharp results with details, while per-frame processing makes the result video temporal inconsistent and jitters across frames. On the contrary, our zero-shot video restoration framework restores a low-quality input video into a temporally consistent high-quality video.
@@
Video denoising, compared to VSR, is a simpler task for regression models, as they can often find the correct pixel value given a sufficiently large batch size. However, our method consistently outperforms others in terms of visual quality (LPIPS) and remains highly robust even as degradation becomes severe. Fig. 7 visualizes the denoising results on the REDS30 dataset. Shift-Net (Li et al., 2023) fails to remove all noise, likely due to the out-of-domain noise level; VidToMe (Li et al., 2024) produces smooth results but lacks fine details. Although DiffBIR (Lin et al., 2024) generates highly detailed images, it suffers from poor temporal consistency, as evident in the changes to the pedestrian’s head and the statue’s face. In contrast, our method preserves both fine details and temporal consistency, effectively balancing these two aspects.
Other Video Tasks: Consistent Video Depth.
Our zero-shot framework is applicable to any pre-trained image-based diffusion models and could improve the predicted video consistency. Therefore, we integrate our proposed zero-shot framework into a state-of-the-art latent diffusion-based monocular depth estimator: Marigold (Ke et al., 2024). Fig. 8 shows that integrating our proposed framework into Marigold helps improve the temporal consistency of video depth estimation. We provide more visual comparisons in the supplementary materials.
| Input | ![]() |
| Marigold | ![]() |
| Ours | ![]() |
This adaptability to various tasks (super-resolution, denoising, depth estimation) showcases the broad applicability of our approach. As more powerful or specialized image models emerge, our framework can quickly adapt to leverage these improvements for video restoration tasks. We provide computational complexity evaluations in the supplementary materials.
@@ 4.2 Ablation Study
Tab. 2 presents an ablation study comparing different approaches (optical flow and cosine similarity) for finding correspondences and their order in the UNet. As detailed in Sec. 3.3, the hybrid approach of using optical flow at the downsample blocks and cosine similarity at the upsample blocks achieves the best performance. Additionally, our proposed spatial-aware token merging further enhances performance by utilizing spatial information to guide correspondences. See supplementary materials for temporal profile comparisons.
@@ 5 Conclusion
Our framework has two main limitations: (1) LDM decoder sensitivity can cause flickering in dynamic scenes. (2) Extreme degradation may yield unsatisfactory results. Future work will address these issues by stabilizing decoder output, and enhancing extreme degradation handling. Our framework’s adaptability allows for the integration of future, more powerful diffusion models.
@@ Appendix A Appendix / supplemental material
A.1 Ablation studies on correspondences identified by cosine similarity
Fig. 9 The figure shows the correspondences at denoising step 40 for three scenarios: without spatial awareness and padding removal, without spatial awareness, and with both spatial awareness and padding removal (ours). It is evident that padding values significantly affect the matching quality. However, even after removing padding, many mismatched diagonal lines remain, leading to blurry results. In contrast, our method effectively finds accurate correspondences by leveraging spatial information from the video.
A.2 Severe Degradation Scenarios.
Our balanced approach proves particularly effective in severe degradation scenarios. For instance, in 8 super-resolution tasks, our method not only avoids artifacts but can even improve visual quality compared to per-frame approaches (Fig. 10). This underscores the effectiveness of our ratio annealing technique in addressing the over-smoothing tendency while maintaining the benefits of our token merging approach. Additional comparisons on video super-resolution can be found at Fig. 11 and Fig. 12.
Other Video Tasks: Consistent Video Depth.
Our zero-shot framework is applicable to any pre-trained image-based diffusion models and could improve the predicted video consistency. Therefore, we integrate our proposed zero-shot framework into a state-of-the-art latent diffusion-based monocular depth estimator: Marigold (Ke et al., 2024). Fig. 13 shows that integrating our proposed framework into Marigold helps improve the temporal consistency of video depth estimation.
| Input | ![]() |
| Marigold | ![]() |
| Ours | ![]() |
| Input | ![]() |
| Marigold | ![]() |
| Ours | ![]() |
| Input | ![]() |
| Marigold | ![]() |
| Ours | ![]() |
| Input | ![]() |
| Marigold | ![]() |
| Ours | ![]() |
A.3 Computational Complexity
While our method focuses on zero-shot video restoration without additional training, it’s important to consider the computational requirements in comparison to other approaches. Tab. 3 provides an overview of the training time and GPU specifications for different methods, including ours.
As shown in the table, our method stands out by not requiring any training or fine-tuning, which significantly reduces the computational resources needed. This is in stark contrast to other methods that require multiple high-end GPUs and several days of training time. For inference, our method introduces some computational overhead due to the hierarchical latent warping and hybrid token merging processes. However, this overhead is relatively small compared to the resources required for training or fine-tuning video models. Specifically, our method adds only approximately 6 seconds to the inference time of the base image diffusion model per frame.
| Method | Training time | GPU specs |
| Shift-Net (Yan et al., 2018) | Not reported | 8 NVIDIA A100-32G GPUs |
| FMA-Net (Youk et al., 2024) | Not reported | Not reported |
| Upscale-A-Video (Zhou et al., 2023) | Not reported | 32 NVIDIA A100-80G GPUs |
| Ours | No training needed | - |
A.4 Additional Ablation Studies
Comparison of temporal profiles.
The comparisons in Fig. 14 also indicate that our results are smoother, demonstrating better temporal stability.
Token Unmerging Strategies.
We experimented with two unmerging strategies: averaging paired tokens and direct replacement with keyframe tokens. Tab. 4 shows the results of these experiments on the Vid4 x4 SR task. As shown in the table, the replacement method outperforms averaging in terms of LPIPS, indicating better perceptual quality. Our experiments consistently showed that averaging tends to produce blurrier outputs in restoration tasks. Based on these results, we adopted the replacement-based unmerging process in our final model, as it preserves more details and leads to sharper outputs.
| Unmerging Method | LPIPS |
| Averaging | 0.337 |
| Replacement | 0.329 |
Limitations: Extreme Degradation
Extreme degradation (e.g., 32 super-resolution) or overly detailed facial features may yield unsatisfactory results (Fig. 15). However, our framework’s adaptability allows the incorporation of future, more powerful image-based diffusion models. Future improvements will focus on refining keyframe selection, stabilizing decoder output across LDM architectures, and enhancing extreme degradation handling. These aim to improve practical application and mitigate flickering issues inherent in LDM decoders.
| (a) DDNM | (b) FMA-Net | (c) Ours |